# DUKAE: DUal-level Knowledge Accumulation and Ensemble for Pre-Trained Model-Based Continual Learning

### v1

@@ 4.2 Decision-level Knowledge Accumulation
$\begin{split}\min_{\varphi_{k}}&\left[\sum_{(x,y)\in D_{t}}\mathcal{L}\left(g(f(x;\theta,\mathcal{P}_{k});\varphi_{k}),y\right)\right.\\ &\left.+\sum_{c=1}^{|C_{1:t-1}|}\sum_{(x_{c},y_{c})\sim\mathcal{G}_{k,c}}\mathcal{L}\left(g(x_{c};\varphi_{k}),y_{c}\right)\right],\end{split}$ (6)
@@
Table 1: Result of LAA and IAA for baseline methods and our approach across four datasets. * means results from our re-implementation.
Method CIFAR-100 ImageNet-R CUB-200 Cars-196
LAA (%) IAA (%) LAA (%) IAA (%) LAA (%) IAA (%) LAA (%) IAA (%)
L2P [32] 82.76±1.17 88.48±0.83 66.49±0.40 72.83±0.56 62.21±1.92 73.83±1.67 38.18±2.33 51.79±4.19
DualPrompt [31] 85.56±0.33 90.33±0.33 68.50±0.52 72.59±0.24 66.00±0.57 77.92±0.50 40.14±2.36 56.74±1.78
CODA-Prompt [27] 86.56±0.77 90.61±0.36 75.25±0.56 81.26±0.76 72.63±0.76 80.54±0.54 44.89±0.61 58.91±0.37
LAE [8] 85.59±0.46 89.96±0.44 72.66±0.63 78.91±0.89 77.48±0.94 85.83±0.68 52.47±1.46 64.08±1.01
ADAM [36] 87.46±0.03 92.20±0.06 66.70±0.21 75.18±0.07 86.77±0.02 91.32±0.01 41.77±7.73 53.78±7.76
EASE [37] 87.73±0.17 92.29±0.19 75.89±0.28 81.67±0.20 86.62±0.08 91.20±0.03 37.58±0.25 51.00±0.14
SLCA [33] 91.53±0.28 94.09±0.87 77.00±0.33 81.17±0.64 84.71±0.40 90.94±0.68 67.73±0.85 76.93±1.21
SLCA++ [34] 91.69±0.15 94.47±0.72 79.78±0.16 84.31±0.73 86.59±0.29 91.63±0.72 73.97±0.22 79.46±0.80
RanPAC [24] 92.10±0.17 95.13±0.06 77.44±0.34 83.06±0.30 89.09±0.23 92.86±0.03 74.69±0.92 82.24±0.41
4.3 Adaptive Expertise Ensemble
@@
$\hat{y}=\arg\max_{y}\left(\text{Concat}\left[S_{1}(x),S_{2}(x),\dots,S_{T}(x)\right]\right).$ (8)
5 Experiments
5.1 Benchmark and Evaluation Metrics
We follow SLCA [33] to evaluate our method on four widely-used benchmarks—CIFAR-100 [16], ImageNet-R [9], CUB-200 [29], and Cars-196 [15]—where the first two datasets focus on relatively coarse-grained classification tasks, while the latter two emphasize fine-grained classification. CIFAR-100 consists of 100 classes of small-scale images across 100 classes, with 50,000 training images and 10,000 test images. ImageNet-R comprises 200 classes of large-scale images, with 24,000 train images and 6,000 test images styled in various artistic renditions. CUB-200 focuses on fine-grained bird classification which contains 200 bird species and split into 5,994 training images and 5,794 test images. Cars-196 dataset includes 16,185 images of 196 fine-grained vehicles which is divided into 8,144 training images and 8,041 test images. In our approach, all datasets are divided into 10 tasks to simulate the continual learning scenario. Each task in CIFAR-100 comprises 10 classes, while in ImageNet-R and CUB-200, each task contains 20 classes. For Cars-196, the first task consists of 16 classes, with each subsequent task including 20 classes.
Regarding the evaluation metrics, we follow [33, 34] to present the average accuracy of all classes after learning the last task, which we refer to as Last task Average Accuracy (LAA, corresponding to “Last-Acc” in [33, 34]). We also calculate the average accuracy of all previously seen classes after learning each incremental task, which we denote as Incremental tasks Average Accuracy (IAA, corresponding to “Inc-Acc” in [33, 34]).
5.2 Baselines
We compare our proposed method against several PTMCL approaches, including L2P [32], DualPrompt [31], CODA-Prompt [27], LAE [8], ADAM [36], EASE [37], SCLA [33], SLCA++ [34] and RanPAC [24]. Due to differences in result presentation and specific experimental settings, we reproduced the results of the ADAM, EASE, and RanPAC methods using the authors’ official open-source code, ensuring the use of the same PTM parameters and maintaining the same class order in the datasets for a fair comparison. For the remaining methods, their experimental results are primarily referenced from SLCA++ [34]. To ensure a fair comparison, we follow [36] to apply a random seed 1993 to shuffle the class order of the dataset before dividing it into 10 tasks.
@@ 5.2 Baselines
Figure 5: Performance comparison of ensemble with varying number of subspaces. Each figure shows LAA results from the first task to the tenth task using varying number of subspaces for ensemble across four datasets. Each line in the figure represents the performance of the ensemble with different number of subspaces after learning the corresponding task. For the first task, a maximum of one subspace can be utilized for ensemble, and this increases incrementally such that the tenth task allows for the use of up to ten subspaces for ensemble.
5.3 Training Details
5.4 Results Analysis
5.5 Ablation Study
Table 2: LAA results with and without SSL loss ($\mathcal{L}_{\text{SSL}}$) applied to learn PEFT modules across four datasets.
@@ 5.5 Ablation Study
About Ensemble Strategy. To validate that our proposed adaptive expertise ensemble can effectively integrate useful information from different subspaces, we compared it with a simple ensemble method, which directly applies averaging of the classification results from all subspaces, and also with the strategy in which no ensemble is used (only using the result of subspace corresponding to the first task). The results are shown in Table 3. As can be seen in the table, the simple ensemble method cannot effectively integrate the knowledge from each feature subspace, and its performance is significantly lower than that of our proposed adaptive expertise ensemble. In fact, in most cases, its performance is even worse than the no-ensemble approach, which indicates that the simple ensemble method fails to properly utilize the knowledge from each subspace and may even be detrimental to performance. This further provides evidence that our proposed adaptive expertise ensemble can integrate and utilize knowledge from each subspace correctly and efficiently.
Table 3: LAA results with no-ensemble (NoE), simple ensemble (SE) and our adapter expertise ensemble (AEE).
@@
Table 4: Results with different continual learning tasks. The first four dataset results (top block) are obtained using Adapter, while the latter four (bottom block) correspond to LoRA.
@@
Figure 6: Adaptive expertise ensemble analysis. Each figure shows accuracy of each task when performing ensemble with varying number of subspaces after completing the 10th task. Each row in the figure represents the accuracy of each task when using an ensemble with varying number of subspaces.
5.6 Adaptive Expertise Ensemble Analysis

### v2

@@ 4.2. Decision-level Knowledge Accumulation
(6) $\begin{split}\min_{\varphi_{k}}&\left[\sum_{(x,y)\in D_{t}}\mathcal{L}\left(g(f(x;\theta,\mathcal{P}_{k});\varphi_{k}),y\right)\right.\\ &\left.+\sum_{c=1}^{|C_{1:t-1}|}\sum_{(x_{c},y_{c})\sim\mathcal{G}_{k,c}}\mathcal{L}\left(g(x_{c};\varphi_{k}),y_{c}\right)\right],\end{split}$
@@
4.3. Adaptive Expertise Ensemble
@@
(8) $\hat{y}=\arg\max_{y}\left(\text{Concat}\left[S_{1}(x),S_{2}(x),\dots,S_{T}(x)\right]\right).$
@@
Figure 5. Performance comparison of ensemble with varying number of subspaces. Each figure shows LAA results from the first task to the tenth task using varying number of subspaces for ensemble across four datasets. Each line in the figure represents the performance of the ensemble with different number of subspaces after learning the corresponding task. For the first task, a maximum of one subspace can be utilized for ensemble, and this increases incrementally such that the tenth task allows for the use of up to ten subspaces for ensemble.
5. Experiments
5.1. Benchmark and Evaluation Metrics
We follow SLCA (Zhang et al., 2023) to evaluate our method on four widely-used benchmarks—CIFAR-100 (Krizhevsky and Hinton, 2009), ImageNet-R (Hendrycks et al., 2021), CUB-200 (Wah et al., 2011), and Cars-196 (Krause et al., 2013)—where the first two datasets focus on relatively coarse-grained classification tasks, while the latter two emphasize fine-grained classification. CIFAR-100 consists of 100 classes of small-scale images across 100 classes, with 50,000 training images and 10,000 test images. ImageNet-R comprises 200 classes of large-scale images, with 24,000 train images and 6,000 test images styled in various artistic renditions. CUB-200 focuses on fine-grained bird classification which contains 200 bird species and split into 5,994 training images and 5,794 test images. Cars-196 dataset includes 16,185 images of 196 fine-grained vehicles which is divided into 8,144 training images and 8,041 test images. In our approach, all datasets are divided into 10 tasks to simulate the continual learning scenario. Each task in CIFAR-100 comprises 10 classes, while in ImageNet-R and CUB-200, each task contains 20 classes. For Cars-196, the first task consists of 16 classes, with each subsequent task including 20 classes.
Regarding the evaluation metrics, we follow (Zhang et al., 2023, 2024a) to present the average accuracy of all classes after learning the last task, which we refer to as Last task Average Accuracy (LAA, corresponding to “Last-Acc” in (Zhang et al., 2023, 2024a)). We also calculate the average accuracy of all previously seen classes after learning each incremental task, which we denote as Incremental tasks Average Accuracy (IAA, corresponding to “Inc-Acc” in (Zhang et al., 2023, 2024a)).
5.2. Baselines
We compare our proposed method against several PTMCL approaches, including L2P (Wang et al., 2022c), DualPrompt (Wang et al., 2022b), CODA-Prompt (Smith et al., 2023), LAE (Gao et al., 2023), ADAM (Zhou et al., 2024a), EASE (Zhou et al., 2024b), SCLA (Zhang et al., 2023), SLCA++ (Zhang et al., 2024a) and RanPAC (McDonnell et al., 2024). Due to differences in result presentation and specific experimental settings, we reproduced the results of the ADAM, EASE, and RanPAC methods using the authors’ official open-source code, ensuring the use of the same PTM parameters and maintaining the same class order in the datasets for a fair comparison. For the remaining methods, their experimental results are primarily referenced from SLCA++ (Zhang et al., 2024a). To ensure a fair comparison, we follow (Zhou et al., 2024a) to apply a random seed 1993 to shuffle the class order of the dataset before dividing it into 10 tasks.
Table 2. LAA results with and without SSL loss ($\mathcal{L}_{\text{SSL}}$) applied to learn PEFT modules across four datasets.
@@ 5.2. Baselines
5.3. Training Details
5.4. Results Analysis
5.5. Ablation Study
Table 3. LAA results with no-ensemble (NoE), simple ensemble (SE) and our adapter expertise ensemble (AEE).
@@ 5.5. Ablation Study
About Ensemble Strategy.
To validate that our proposed adaptive expertise ensemble can effectively integrate useful information from different subspaces, we compared it with a simple ensemble method, which directly applies averaging of the classification results from all subspaces, and also with the strategy in which no ensemble is used (only using the result of subspace corresponding to the first task). The results are shown in Table 3. As can be seen in the table, the simple ensemble method cannot effectively integrate the knowledge from each feature subspace, and its performance is significantly lower than that of our proposed adaptive expertise ensemble. In fact, in most cases, its performance is even worse than the no-ensemble approach, which indicates that the simple ensemble method fails to properly utilize the knowledge from each subspace and may even be detrimental to performance. This further provides evidence that our proposed adaptive expertise ensemble can integrate and utilize knowledge from each subspace correctly and efficiently.
Table 4. Results with different continual learning tasks. The first four dataset results (top block) are obtained using Adapter, while the latter four (bottom block) correspond to LoRA.
@@
Figure 6. Adaptive expertise ensemble analysis. Each figure shows accuracy of each task when performing ensemble with varying number of subspaces after completing the 10th task. Each row in the figure represents the accuracy of each task when using an ensemble with varying number of subspaces.
5.6. Adaptive Expertise Ensemble Analysis